August 20, 2020 - ME/INJ/CIS/CASA

From Ciswikidb
Jump to: navigation, search

Meeting Info

We will meet Thursday @ 10am EDT via BlueJeans
Meeting URL (for video and/or audio call in):
Phone number (for audio call in): 1-888-240-2560
Meeting ID: 909497334#

Meeting August 20

  • SAD begins Sep 21
Region 1
Region 2
Region 3
Region 4

Meeting August 13

Region 1
  • Marcy vacuum bake 400 C gun/y-chamber
  • Are there (4) holes in the pedestal?
Region 2
  • (10) tophats + (10 plates)
  • Finish solenoid 2 soon
  • Add supports/holes to support prebuncher if needed
  • Lakshmi would like Wien quads further out radially by 3.2cm, she'll run worst/best case models, meeting next Tue/Wed
  • Carlos wrapping up UITF Wien tests, likely will ask for (2) new pairs of electrodes
Region 3
  • MHD coil design completed, waiting on bids
Danny updates
  • 3 wien filter fixtures ordered.
  • some spool pieces ordered.
  • 5 fixtures for MFX solenoids waiting for procurement to order.
  • haven't seen the NEG flange yet, supposedly shipped monday.
  • magnet vendors working on MHD quotes, waiting for final drawings.
  • both bunchers in machine shop for first trimming after RF tuning. should be done in a couple of days.
  • submitted order to the machine shop for modifying A1 and A2 cubes.

Meeting August 6



  • First solenoid in TL 1137. The magnet map starts with the side away from the water lines. The first lead is labelled 1. Phil has confirmed orientation and drawing for rotatable and non-rotatable flanges.
  • Second solenoid in magnet lab for mapping, finishing installing on v-blocks, mapping to start soon
  • The next solenoids are in process, hope to see third solenoid next week?

15 deg Dipole magnet update: Mike B. They have worked on a dummy coil, first production coil has been started.

Aperture correctors: drawing done, Danny has out to the usual suspects for quote process.

 *Boltup on top of cubes will require welding a fixture to the cube since there is no room for bolts adjacent to the aperture manipulators and magnets. 
 *Cubes also need machining for survey tooling balls
 *Add labels – engraved or punched – to the cubes indicating orientation
 *Danny working on print

From Lakshmi:

 *MHD integrated field homogeneity is 99.91 % at 2 mm radius
 *The structural evaluation of the MHD coil assembly is in progress; expected to have the results this evening.
Vacuum hardware

Y chamber girder

 *Bottom flange on track to arrive next week
 *garages for laser and y chamber viewers are ready to bend at the right length and tack weld

Things ordered

 *NEG pumps for apertures just sent order ing
 *NEG pumps and supplies for wien filters and laser chamber have arrived
 *vertical flipper spool has been ordered
 *Adapter spools for aperture cubes still being finalized.

Buncher has been brazed, see detail below

Wien filter

 UITF assembled with ceramic bolts and washers to avoid high gradients on bolt head
   Consideration of very mild spring washer in ceramic bolt assembly
   Electrode plate will expand during bake, Ceramic bolt won’t expand,could strip Macor bushing
   Very thin spring washer could be used to protect macor bushing during bakeout from expansion and stripping. 
   Danny and Carlos will continue discussion

Region 1
Region 2
Region 3
  • Buncher cavity fabrication update
The ports have been brazed in and are leak tight. (three ports on the body, two end tubes) the machine shop removed some tooling that was accidentally brazed to the i.d. of one of the ports and rounded out one of the cavity bodies that had gotten out-of-round due to brazing.  Haipeng has clamped everything together and done his first test and trimming calculation. Haipeng will take both cavities to the shop today and have them bore out the bodies to 1/2 of what the's calculated the trimming to be.
Region 4

Meeting July 30

Region 1
  • Small shed electrode has been surveyed, back in the lab, needs CO2 jet and store
  • First round of NEG coating light, so Phil will continue, wrap up next week.
  • Musson is ready to map BPM's as soon as ready.
  • Mike finishing FX-1 mapping, rocking and rolling.
Region 2
  • Danny says Norcal flange due to ship Aug 10
  • Phil needs measurement from Shaun for correction insertion, then will get them cut and tacked welded chamber (Phil has 5)
  • McMaster Carr stand-offs, the oldest one
  • From Milhous: We wound a “dummy” coil and tweaked the tooling. The first production coil is a couple layers into winding and looks good. We probably need two weeks to finish both coils.
Region 3
  • Bubba re-assembled the UITF Wien filter w/ ceramic screws an standoffs (instead of metal)
  • Wien bake 160-175 (Carlos concerned about the bake-out) Carlos concerned about breaking Ceramic (needs spring washer)
  • Wien goes to UITF next Tuesday, and high voltage testing when time there is no beam
  • Wien fixture design completed, Danny starting to send for bids (to get 3)
  • Danny thinks buncher jobs are moving OK w/o additional meeting for it
  • Drawings for pedestals needs to be completed, but will happen after A1
  • Last week Scott Williams brazed one part OK, another part came out of round and a tool part stuck, machine shopped fixed found
  • Haipeng has one cavity, will start testing, then take to machine shop for machining
  • Danny is asking Haipeng to double-check the numbers
Region 4
  • Lakshmi presented new HD coil design, will use MWS 14 AWG copper wire, has 50% overhead (design ~2A, operate to ~3A), quantifying next order six-pole within r=2mm, can use 3A or 10A card, will need a clamp shell for the heat treatment of 200-250 C.
  • Joe G. will contact Sarin about trim card, up to 3.4 A, copy Lakshmi
  • Gary almost done w/ detail of HD bobbin, will then send to Lakshmi for mechanical analysis
  • Shaun and Gary are finishing the A1/A2 design, and spoke to Chris Gould about tooling ball location
  • Joe started listing names for song sheet, will circulated for review, and then to Danny for songsheet update, then distributed to installation teams

Week of July 16

Region 1
  • Electrode does not fit on to the ceramic, Bubba will grind/sand the ceramic down expect about half-mil
  • Sputtering
  • 7/16/2020 - Update
::MFX Solenoids (ETI)
::1) MFX #001 is on the QC table getting electrical and flow testing done to confirm acceptance.  Magnet measurement should start next week.
::2) MFX #002 is en route to the Lab.  ETA Friday.
::3) Completed winding coil #12  so half way done with all coils.
  • Mike is going to recommend LCW line for solenoid, Shaun will check if it fits at all locations
Region 2
  • NEG flange due early August
  • Shields need to be larger dia
  • MDS 15 degree bend Dipoles (Milhous) Finally in contact with POC Joel Meissner. Winding is occurring now on first coil. They lost two technicians in their shop and had to hire/train new help. Our coils are being wound by another technician who had to be pulled off another job. He promises to follow up with pictures next week.
Region 3
  • Bubba submitted work order for 30 longer stand-offs
  • Danny looking at dwgs for 11 and 13 mm gap springs, finish design, test Wien 15 mm gap before buying smaller gap plates
  • Scott brazing prebuncher, then SRF leak testing
Region 4
  • Gary working on A1/A2 supports
  • When can Lakshmi return to the MHD and Wien quads

Week of July 9

=> Dennis, Alicia, Yan, Joe work together to pick names, then Dennis will populate Elegant workspace/CED, Danny will make songsheet, we will celebrate!!!

Region 1
  • Small shed back from the shop, Bubba polishing tee/rear at home, then to S&A
  • First article solenoid to be QC's, second solenoid will shipped, Joe send mapping info, Phil give solenoid
  • Strung NEG wire Tuesday, prepping for characterization/thickness, bake then coating starting Friday...
  • Shaun says Musson's stand parts are on his desk + in shop for pick-up
Region 2
  • Joe reached out several times to dipole machine shop (June 25th), no word from them in a few weeks
  • Shaun will talk to vendor on viewer shields
  • Shaun finishing viewer top hats for Danny review
  • Danny says NEG flange due Aug 6
Region 3
  • Mike Beck finishing UITF blue wien thermal/B measurements to 20 A (went to 15 A only so far)
  • Danny, find long macor drawing (need 24 long)
  • Buncher meeting 1:30
Region 4

From Danny, in today's meeting, we should address these items:

  • A1...
did you want to adjust the extended and retracted positions of A1? see sketches on email sent 7/2. media:200702_old A1 vs new A1.pdf => Yes
did we want to make alignment fixtures for A1 and A2 cubes? => Yes
do we want a bellows downstream of the DP can? => Yes
Agreed at least > 1/16th...ID clearance for MHD coils
  • A2...
do we want to re-use what we have there? Yes, also make a spare? Yes, do we already have spares? is there really an A1 modified there. Yes

Week of July 2

Week of June 25

  • Mike finding small shorts to ground on Wien coils
  • Solenoid first article arrived
  • Joe M. schedule meeting for field maps and probs with Wien coils
  • Carlos sends Mike assembly drawing
Region 1
  • Last antennae arrived at shipping/receiving
  • Bake of parts going well, ends next week, then NEG coating
  • Phil/Marcy have ceramics/NEG wire
Region 2
  • SG sent to get (3) garage shields fabricated w/ Vince
  • SG still working on the tophats, designed a fixture w/ same cross-
  • Bottom Y-flange ordered (heat treat)
  • Have all pumps, vacuum parts
  • JG send Danny info on the laser vacuum window
  • MS/PA talk w/ Tony about whether to redo bellows shaft/turnbuckle mechanism
Region 3
  • Joe met w/ Lakshmi, she'll start evaluating a strong QW mid-July
  • Brock done w/ copper cavities, shipping to JLab ~next, Matt syncing up with Musson
Region 4
  • Existing HD coils likely okay, Reza will confirm; Lakshmi will model coils to verify assumptions
  • Shaun/Gary want to order cube flanges; we'll pull trigger once HD design fixed
  • Danny is working on the aperture design

Week of June 11

  • brackets arrived (10) 2-3/4 and (5) 4-/12
Region 1
Region 2
  • Assembly drawing for Y-girder not yet complete
  • SG checking top-hat dwgs w/ DM
  • SG give machine shop dwg of tube to check fit-up
  • Milhous says fixtures done, coils going winding machine today
  • Proc should make order of NEG flange today (8 week delivery time)
  • Missing 2nd order (vee-block + fixture plate)
  • 2nd solenoid
  • Have 2nd yellow spool
Region 3

Last week June 4


MFX Solenoids (ETI)
  • Waiting on update on the measurements of the 1st solenoid.
  • The 2nd solenoid coils completed and passed all electrical testing.
  • In process of winding coils for the 3rd solenoid.
15 deg. bend dipole (Milhous)
  • Coil tooling is ready.
  • Coil winding will proceed week of JUN8 (when winding machine becomes available).
  • Discussing their QA process with the coils.
  • Gathered documentation on the magnet from the PR and emails.

We'll go back to work next week....

Last week May 28 talking points

Plan for the upcoming SAD and injector rebuild (Grames, Adderley, Marcy, Carlos, , Bubba, Windham)

  • a. Long list of jobs for new NEG BPM's here:
  • b. Capture new solenoid on NEG pipe, NEG coat them too
  • c. Assemble new Y-chamber on girder, new bottom flange, new NEG's, finish coating that chamber
  • d. Setup vacuum chamber if Caryn is going to evaluate any windows
  • e. First or second, provide UITF Wien magnet to Meyers (Phil), hi-pot and map 3 coils pairs
  • f. Second or first, rebuild UITF Wien internals to eliminate high E-points (Bubba), finalize spacers/springs/gap (Carlos/Danny remotely)
  • g. Setup rebuilt Wien in UITF and HV test (Carlos)
  • h. Work w/ Shaun to test top-hats and garage needed for John's y-chamber design
  • i. Collect up dozens of vacuum parts, clean, and get ready for SAD (pipes, bellows, top-hats)
  • j. Work w/ Danny/Shaun/S&A to fabricate and assemble new A2 aperture on a cube
  • k. Work w/ Scott to assembly dummy of motorized A1 on new cube, ready for shutdown
  • l. 11 Collect up all the girders, supports, hardware
  • m. Build two prebuncher cavities?
  • n. Work w/ shop to capture Wien quads on crosses, weld on flanges, for tunnel viewers
  • o. Build up and vacuum test the y chamber: RGA, base pressure
  • p. Build up and test the docking chamber as needed (won’t be needed until later in the shutdown)

Last Week's Talking Points - May 21

  • Progress!
First coat of green paint on "core"
Y/Laser chamber viewers
  • Design status
Wien Filter
From Mike Beck - The 6 new "Wien" coils meet dimensional check acceptance criteria and counted turn count all as required at 2 layers x 13 turns per. Will put spread sheet with dimensional data together in coming days. Also, checked resistance on all 6, the range was consistent at ~.2 ohms. The visual quality of coils that Technicoil provided are nice in appearance.  Will conduct hipot testing when able to get back onsite and confirm the rest of QA acceptance testing, fit to core (when core arrives), and mapping required.
Wien coils in crate at MMF
MHD A1/A2 Steering Coils
PSS Kicker
  • ME model status
  • Getting quotes (2-4)
Wien quads
  • CED from Gary Croke => MQW: B'L = 15 G/A with +/- 3A limits
  • Will circulate when done